A ball is hit with an initial velocity of 15 m/s and follows a parabolic motion path. How long does it take for the ball to reach it's maximum height?
Answer #1

soi:- niven : u=15 mis As we know Vf=U-at v=15 -9,8 xt (At mox height veco) 0315-987 5=9:8€ It1:53)